Kaden is 1/4 as old as Tom now. In 7 years, the sum of their age will be 54. How old is Tom now?
Answer by mananth(16946)

Kaden is 1/4 as old as Tom now. In 7 years, the sum of their age will be 54. How old is Tom now?
Kaden(K) is 1/4 as old as Tom(T) now.
k=T/4
In 7 years, the sum of their age will be 54.
K+7+T+7=54
K+T=40
substitute k
T/4 +T =40
5T/4 = 40
5T= 160
T= 32 years